# Geminga: A Window of the Role Played by Local Halo in the Cosmic Ray Propagation Process

### v1

@@ Abstract
An emerging commonality among the recently observed pulsar halos is the presence of distinct radiation patterns at high energies, while no extended radiation is detected around the GeV energy band. This commonality suggests that pulsar halos play a crucial role in the local propagation of cosmic rays, making it necessary to investigate the underlying mechanisms of this phenomenon. This work focuses on the 3D propagation study of cosmic rays, incorporating the Geminga pulsar into our propagation model to investigate its contribution to different observational spectra. We consider Geminga a dominant local source of positrons, successfully reproducing the observed positron spectrum and multi-energy spectra of the Geminga halo. Through calculations of signal and background at different angles, we find that: (1) at low energies, the slow diffusion characteristic around the pulsar region leads to a low electron density in the extended area around Geminga, causing the background radiation to exceed the signal intensity far; (2) at high energies, the larger effective diffusion radius of high-energy electrons/positrons causes the signal from Geminga to dominate the local high-energy phenomena; (3) the observed fluctuation of diffuse gamma-ray radiation by LHAASO is likely due to the incomplete subtraction of radiation from the local halo. We hope LHAASO will detect more cosmic ray halo sources to validate our model further.
@@ 1 Introduction
As cosmic ray observation experiments enter an era of precision, traditional cosmic ray propagation models cannot predict some finer spectral structures. For instance, primary CR spectra exhibit a spectral hardening at high energies. Data from AMS-02 reveals that the CR spectra above several hundred GeV deviate from a simple power law, displaying a spectral hardening, a phenomenon now confirmed across various experimental platforms (Aguilar et al., 2021; An et al., 2019; Atkin et al., 2018). Secondary CR spectra also harden at this energy range, and their spectra are steeper than those of primary CRs beyond this energy. The latest observations by DAMPE further confirm an excess in the secondary-to-primary flux ratio of CRs at high energies (Dampe Collaboration, 2022).
@@
To explain these observations, a spatially dependent CR propagation model has been developed. In contrast to conventional models that assume uniform propagation of CRs across the Galaxy, we propose the following: (1) CR propagation depends on the local distribution of matter (Guo & Yuan, 2018; Guo et al., 2016; Evoli et al., 2012); (2) local CR sources act as both a potential well, confining local CRs, and a barrier, impeding external CRs from penetrating these regions (Yao et al., 2024; Nie et al., 2024; Tomassetti, 2012); and (3) local CRs primarily consist of two components—component A, comprising CRs accelerated from distant sources and diffused into the region, which dominates at low energies, and component B, consisting of CRs from local sources, which dominate at high energies (Yao et al., 2024; Nie et al., 2024). High-energy phenomena in the Galactic vicinity are thus predominantly driven by local sources. This propagation model effectively accounts for the observed CR spectral features. However, observations of diffuse gamma-ray emissions in the Galaxy and extended emissions from local CR sources may provide further validation of this propagation framework.
@@
Gamma-ray extended emissions observed near pulsars confirm the existence of regions around them that exhibit slower diffusion relative to the interstellar medium (ISM). Since the morphology of pulsar halos follows the spatial distribution of their parent electrons/positrons, they serve as ideal probes for studying local CR propagation in the Galaxy. Observations of extended radiation from pulsars at different energy bands provide an opportunity to test our model. Recently, the Milagro Collaboration (Abdo et al., 2009), High-Altitude Water Cherenkov Observatory (HAWC) (Abeysekara et al., 2017), HESS (H. E. S. S. Collaboration et al., 2023). and Fermi-LAT (Di Mauro et al., 2019) have all revealed radiation from the Geminga halo. Notably, the Milagro Collaboration reported gamma-ray radiation from the Geminga direction extending over 2 degrees and spanning the energy range of 1-100 TeV. This observation was confirmed by HAWC, which detected extended gamma-ray emissions coincident with the position of Geminga. However, a decade-long analysis of Fermi-LAT data indicates no significant extended emission beyond the Geminga region in the 10-500 GeV energy range (Quattrociocchi et al., 2019; Xi et al., 2019). Therefore, we believe that local CR halos play a crucial role in the propagation of CRs in the Galaxy, and investigating the mechanisms that generate the observability of pulsar extended emissions at different energy bands is essential.
@@
This work focuses on the observations of radiation spectra from different energy bands in the Geminga halo, made by Fermi-LAT and HAWC, as well as the distribution of diffuse gamma rays near Geminga observed by LHAASO, and the measurement of cosmic-ray positrons by AMS-02. These observations constrain the injection and diffusion of cosmic-ray particles in the Geminga halo and allow us to investigate the observability of extended radiation from the Geminga halo across different energy ranges. We aim to test the proposed spatially dependent cosmic-ray propagation model. The structure of this paper is as follows: Section 2 introduces the diffusion of electrons/positrons in the pulsar halo; Section 3 discusses our results; and Section 4 provides a summary and outlook for future work.
@@ 2 Diffusion in Halo
The local halo plays an essential role in the cosmic ray propagation process. The fluctuating structure observed in the radial distribution of diffuse gamma rays by LHAASO may be due to the contribution from local sources that cannot be fully subtracted, particularly in the regions near the Geminga and Cygnus bubbles. Therefore, first, we use the spatially dependent propagation model proposed in previous work (see Appendix A) to simulate the distribution of Galactic background cosmic rays and the corresponding background diffuse gamma-ray radiation they produce. Then, we consider the halo as a slow diffusion region and, using the Geminga halo as an example, calculate its radiation spectrum to explain the latest observational data from LHAASO, Fermi-LAT, and HAWC. Finally, we discuss the observability of the Geminga halo signal under the contamination of the background radiation.
The diffusion coefficient around the pulsar, estimated from gamma-ray observations, suggests that it may be several hundred times smaller than the average value across the entire Milky Way. This slow diffusion could be caused by electrons/positrons escaping from the pulsar itself or by turbulent regions generated by cosmic ray particles escaping from the pulsar’s progenitor supernova remnant. Therefore, in our current work, we treat the diffusion of cosmic ray particles from the pulsar halo as a spherically symmetric “two-zone” diffusion model, which consists of slow diffusion within the halo and faster diffusion in the interstellar medium. The diffusion coefficient is mathematically described as (Jóhannesson et al., 2019):
where represents the particle velocity in units of the speed of light, is the normalization of the diffusion coefficient in the general ISM, is the normalization for the diffusion coefficient within the SDZ with radius , is the particle rigidity, and is the normalization (reference) rigidity. The zone between and is a transition layer where the normalization of the diffusion coefficient increases exponentially from to the ISM value .
@@
Here, is break energy, , and are spectral indices for low and high energy bands, and represents the smoothness parameter. In this work, we fixed , , and .
@@ 3 Results
In this section, we present our computational results, which include the multi-band non-thermal radiation energy spectrum of the Geminga halo, the distribution of diffuse gamma rays in the Galactic disk, and the map distribution of signal intensities for the background and halo in the Geminga region. We also discuss the mechanisms that lead to these distributions and shapes forming near the pulsar. In calculating the background signals of the diffuse gamma rays from the Galactic disk, the Geminga halo, and the spectral components of the background radiation in the halo region, it is necessary to compute the distribution of cosmic ray particles in the Milky Way. We use the B/C and cosmic ray proton spectra observed near Earth to constrain the relevant parameters of the propagation model. Appendix A shows that we use a spatially dependent cosmic ray propagation model to accurately reproduce the observed B/C and proton spectra, with the values of the relevant parameters listed in Table 1.
@@ 3.1 Emissions from the Geminga pulsar halo
To reproduce the observed phenomena from the Geminga halo, we used the model constructed using the method described in Section 2 to calculate the radiation energy spectrum and its morphological distribution from the Geminga halo. Figure 1 presents the multi-wavelength non-thermal radiation from the Geminga halo. Following the observational data, the synchrotron component is limited to the contribution within 0.2 degrees around the pulsar. In contrast, the high-energy IC component comes from the contribution of electron/positron pairs interacting with the cosmic background radiation via inverse Compton scattering in the 10 region. By using the upper limit data from the low-energy segment from NuSTAR and the observational data from Fermi-LAT and HAWC in the high-energy range, the magnetic field of the halo was determined to be . The scale of the pulsar halo’s slow diffusion region is about 50 pc, with a diffusion coefficient . In Figure 2, we present the distribution of the diffuse gamma rays in the Galactic disk. To compare with the LHAASO observational data, we processed the model-calculated data like the LHAASO data: 1) smoothing the theoretical data and 2) subtracting the contributions from point sources and extended sources when calculating the background radiation. Therefore, in this work, the sources detected by KM2A and those observed by other experiments summarized in TeVCat were all subtracted during the calculation of the diffuse radiation, especially in the regions of the Cygnus bubble and Geminga. The region of the Cygnus bubble was masked within about 6, while the region around Geminga had its contribution removed within an 8 radius, which is about 2.5 times larger than the masked regions of other sources listed in TeVCat. As shown in the Figure 2, even with the removal of contributions from extended sources in a larger area, the Cygnus bubble and Geminga halo still provide dominant contributions. The detailed calculation process of the Cygnus bubble component in Figure 2 can be found in our previous work (Nie et al., 2024). Similar to the diffusion of electrons/positrons in the halo in this work, the Cygnus bubble is also considered a slow diffusion region, and its energy spectrum, morphological distribution, and the diffuse gamma-ray structure in the region have been well reproduced.
@@ 3.2 Map of Geminga halo
In the vicinity of the Geminga pulsar, no radiation morphology was observed in the GeV energy range, but there is an extended radiation in the TeV range. To investigate the mechanism behind this phenomenon, we used the observational data of the multi-wavelength energy spectrum from the Geminga halo, as well as the model parameters derived from the constraints of the diffuse gamma-ray observations of the Galactic disk by LHAASO, to calculate the map signals of the Geminga halo and the background. As shown in Figure 3, in the 1-10 GeV energy range, the Geminga halo signal is relatively weak compared to the background intensity, to the point where the extended radiation morphology near the Geminga pulsar is almost invisible due to the contamination of the background. However, in the 1-100 TeV energy range, Geminga produces a signal much stronger than the background, and a significant radiation morphology remains even with the background contribution considered, consistent with the observations from Fermi and HAWC.
@@
Theoretical models suggest that lower-energy electrons/positrons produce the GeV radiation through inverse Compton scattering with soft background photons. However, once the electrons/positrons escape from the pulsar, they propagate through the slow diffusion region near it and diffuse into the more distant and homogeneous normal Galactic environment. In the region where ( ), the diffusion coefficient of particles is at least two orders of magnitude smaller than in other regions due to the spatial-dependence of , meaning that the density of electrons/positrons that generate GeV radiation in this region is low. As a result, the GeV IC radiation flux decreases, and Fermi-LAT can almost not detect any radiation beyond the background.
On the other hand, for the parent electrons/positrons that generate TeV radiation, the larger diffusion coefficient allows them to have a larger effective diffusion radius, enabling them to escape from the slow diffusion region and generate radiation through inverse Compton scattering in a larger extended area. Therefore, there is a strong extended source signal near the Geminga pulsar. This explains why the “bubble” structure remains in the diffuse gamma-ray distribution observed by LHAASO, even after subtracting the contribution from the 8 region around Geminga.

### v2

@@ Abstract
A novel phenomenon among the recently observed Geminga pulsar halo is the presence of distinct radiation morphology at high energies, while no extended radiation is detected in the 10-500 GeV energy band within a region. This phenomenon suggests that pulsar halos play a crucial role in the local propagation of cosmic rays, making it necessary to investigate the underlying mechanisms of this phenomenon. This work focuses on the 3D propagation study of cosmic rays, incorporating the Geminga pulsar into our propagation framework to investigate its contribution to different observational spectra. We consider Geminga a dominant local source of positrons, partially reproducing the observed positron spectrum and multi-wavelength radiative spectra of the Geminga halo. Through calculations of signal and background at different angles, we find that: (1) The slow-diffusion properties near the Geminga pulsar and its proper motion may cause the radiation from electrons originating from Geminga to be distributed across a more extended region. (2) The incomplete subtraction of radiation from the local halo may contribute, to some extent, to the diffuse gamma ray fluctuations detected by LHAASO. We hope that LHAASO will detect more sources of cosmic ray halo to further validate our model.
@@ 1 Introduction
As cosmic ray observation experiments enter an era of precision, traditional cosmic ray propagation models cannot predict some finer spectral structures. For instance, primary CR spectra exhibit a spectral hardening at high energies. Data from AMS-02 reveals that the CR spectra above several hundred GeV deviate from a simple power law, displaying a spectral hardening, a phenomenon now confirmed across various experimental platforms (Aguilar et al., 2021; An et al., 2019; Atkin et al., 2018). Secondary CR spectra also harden in this energy range, and their spectra are steeper than those of primary CRs beyond this energy. The latest observations by DAMPE further confirm an excess in the secondary-to-primary flux ratio of CRs at high energies (Dampe Collaboration, 2022).
@@
To explain these observations, a spatially dependent CR propagation model has been developed. In contrast to conventional models that assume uniform propagation of CRs throughout the Galaxy, we propose the following: (1) CR propagation depends on the local distribution of matter (Guo & Yuan, 2018; Guo et al., 2016; Evoli et al., 2012); (2) local CR sources act as both a potential well, confining local CRs, and a barrier, impeding external CRs from penetrating these regions (Yao et al., 2024; Nie et al., 2024; Tomassetti, 2012); and (3) local CRs primarily consist of two components—component A, comprising CRs accelerated from distant sources and diffused into the region, which dominates at low energies, and component B, consisting of CRs from local sources, which dominate at high energies (Yao et al., 2024; Nie et al., 2024). High-energy phenomena in the Galactic vicinity are thus predominantly driven by local sources. This propagation model effectively accounts for the observed CR spectral features. Besides, observations of diffuse gamma-ray emissions in the Galaxy and extended emissions from local CR sources may provide further validation of this propagation framework.
@@
Gamma-ray extended emissions observed near pulsars confirm the existence of nearby regions with slower diffusion relative to the interstellar medium (ISM). Since the morphology of pulsar halos follows the spatial distribution of their parent electrons/positrons, they serve as ideal probes for studying local CR propagation in the Galaxy. Observations of extended radiation from pulsars at different energy bands provide an opportunity to test our model. Recently, the Milagro Collaboration (Abdo et al., 2009), High-Altitude Water Cherenkov Observatory (HAWC) (Abeysekara et al., 2017), HESS (H. E. S. S. Collaboration et al., 2023), and Fermi-LAT (Di Mauro et al., 2019) have all revealed radiation from the Geminga halo. Notably, the Milagro Collaboration reported gamma-ray radiation from the Geminga direction extending over 2 degrees and spanning the energy range of 1-100 TeV. This observation was confirmed by HAWC, which detected extended gamma-ray emissions coincident with the position of Geminga. However, a decade-long analysis of Fermi-LAT data indicates no significant extended emission around the Geminga region in the 10-500 GeV energy range within a region (Quattrociocchi et al., 2019; Xi et al., 2019). Therefore, we believe that local CR halos play a crucial role in the propagation of CRs in the Galaxy, and investigating the mechanisms that generate the observability of pulsar extended emissions at different energy bands is essential.
@@
This work focuses on the observations of radiation spectra in different energy bands from the Geminga halo, detected by Fermi-LAT and HAWC, as well as the distribution of diffuse gamma rays near Geminga observed by LHAASO, and the measurement of cosmic-ray positrons by AMS-02. These observations constrain the injection and diffusion of cosmic-ray particles in the Geminga halo and allow us to investigate the properties of extended radiation from the Geminga halo in different energy ranges. We aim to test the proposed spatially dependent cosmic-ray propagation model and physical framework. The structure of this paper is as follows: Section 2 introduces the diffusion of electrons/positrons in the pulsar halo; Section 3 discusses our results, and Section 4 provides a summary and outlook for future work.
@@ 2 Diffusion in Halo
The local halo plays an essential role in the cosmic ray propagation process. The fluctuating structure observed in the radial distribution of diffuse gamma rays by LHAASO may be due to the contribution from local sources that cannot be fully subtracted, particularly in the regions near the Geminga and Cygnus bubbles (Nie et al., 2024). Therefore, we first use the spatially dependent propagation model proposed in previous work (see Appendix A) to simulate the distribution of Galactic background cosmic rays and the corresponding background diffuse gamma-ray radiation that they produce. Then, we consider the halo as a slow diffusion region and, calculate its radiation spectrum to explain the latest observational data from LHAASO, Fermi-LAT, and HAWC. Finally, we discuss the properties of the Geminga halo signal under the noise of the background radiation and the proper motion of Geminga pulsar.
The propagation equation for electrons in the Geminga halo is the same as the propagation equation 3 for background cosmic rays. The Geminga halo is treated as a individual cosmic ray source within the Galprop framework. Since diffusion in the halo differs from that in the general ISM environment, diffusion within the Geminga halo is considered to occur in a slow diffusion region. Once the electrons escape from the Geminga halo, their diffusion follows the same scenario as the background cosmic rays. the electrons from Geminga are continuously injected halo around it, and the injection spectrum is normalized so that the total power injected is given by the expression(Jóhannesson et al., 2019) . Here, is the initial spin-down power of the pulsar and , the initial spin-down power is calculated using the current spin-down power of assuming that the pulsar age is .
The diffusion coefficient around the pulsar, estimated from gamma-ray observations, suggests that it may be several hundred times smaller than the average value across the entire Milky Way. This slow diffusion could be caused by electrons/positrons escaping from the pulsar itself or by turbulent regions generated by cosmic ray particles escaping from the pulsar’s progenitor supernova remnant. Therefore, in our current work, we treat the diffusion of cosmic ray particles from the pulsar halo as a spherically symmetric “two-zone” diffusion model, similar to the scenario provided in Fang et al. (2018), which consists of slow diffusion within the halo and faster diffusion in the interstellar medium. The diffusion coefficient is mathematically described as (Jóhannesson et al., 2019):
where represents the particle velocity in units of the speed of light, is the normalization of the diffusion coefficient in the general ISM, is the normalization for the diffusion coefficient within the SDZ with radius , is the particle rigidity, and is the normalization (reference) rigidity. The zone between and is a transition layer where the normalization of the diffusion coefficient increases exponentially from to the ISM value . For the slow-diffusion region of the Geminga halo, we fix the diffusion coefficient index to 1/3. The diffusion coefficient of electrons in the halo is described by Equation 1, while after they enter the ISM, their diffusion coefficient is described by Equation 4 in the Appendix A.
@@
Here, is break energy, , and are spectral indices for low and high energy bands, and represents the smoothness parameter. In this work, we fixed , , and . We have adopted (Schroer et al., 2023), where represents the current spin-down luminosity of the Geminga pulsar in units of .
@@ 3 Results
In this section, we present our computational results, which include the multi-band non-thermal radiation energy spectrum of the Geminga halo, the distribution of diffuse gamma rays in the Galactic disk, and the map distribution of signal intensities for the background and halo in the Geminga region. We also discuss the mechanisms that lead to these distributions and shapes forming near the pulsar. In calculating the background signals of the diffuse gamma rays from the Galactic disk, the Geminga halo, and the spectral components of the background radiation in the halo region, it is necessary to compute the distribution of cosmic ray particles in the Milky Way. We use the B/C ratio and cosmic ray proton spectra observed near Earth to constrain the relevant parameters of the propagation model. Appendix A shows the spatially dependent cosmic ray propagation model used to accurately reproduce the observed B/C and proton spectra, with the values of the relevant parameters listed in Table 1.
@@ 3.1 Emissions from the Geminga pulsar halo
To reproduce the observed phenomena from the Geminga halo, we use the model constructed using the method described in Section 2 to calculate the radiation energy spectrum and its morphological distribution from the Geminga halo.
Figure 1 and 2 present the multi-wavelength non-thermal radiation and the angular profile of the surface brightness from the Geminga halo, respectively. Following the observational data, the synchrotron component is limited to the contribution within 0.2 degrees around the pulsar. In contrast, the high-energy IC component comes from the contribution of electron/positron pairs interacting with the cosmic background radiation via inverse Compton scattering in the 10 region. By using the upper limit data from the X-ray upper limits (XMM–Newton, NuSTAR) and the observational data from Fermi-LAT and HAWC in the high-energy range, the magnetic field of the halo was required to be , which is significantly weaker than the typical magnetic field in the interstellar medium. As shown in Figure 1, It is evident that the theoretical prediction for does not align with the XMM-Newton upper limits, suggesting it may be less effective in constraining the model parameters compared to utilizing the SED data. The other constrained quantity, the scale of the pulsar halo’s slow diffusion region is about 50 pc, with a diffusion coefficient .
Figure 3 shows the emissions produced by the electrons and positrons within the different extended regions around the Geminga pulsar. It can be concluded that low-energy radiation primarily originates from the extended regions farther away from the pulsar, while the high-energy component is dominated by contributions from electrons in the regions closer to the pulsar. This energy-dependent scenario arises because cooling effect of cosmic ray electrons, low-energy electrons have a longer propagation distance compared to high-energy electrons.
@@
In Figure 4, we present the distribution of the diffuse gamma rays in the Galactic disk. The background component of Figure 4 represents the radial distribution of diffuse gamma-ray radiation produced by the interaction of background cosmic rays with ISM gas in the Galactic disk(). We calculate the spatial distribution of cosmic rays using the spatially dependent propagation model described in the Appendix A and utilize the GALPROP package(Strong & Moskalenko, 1998; Porter et al., 2022) to compute the Galactic diffuse gamma-ray radiation, further simulating the radial distribution of diffuse gamma-ray radiation in the Galactic disk. On the other hand, to compare with the LHAASO observational data, we process the model-calculated data like the LHAASO data: 1) smoothing the theoretical data and 2) subtracting the contributions from point sources and extended sources when calculating the background radiation. Therefore, in this work, the sources detected by KM2A and those observed by other experiments summarized in TeVCat are all subtracted during the calculation of the diffuse radiation, especially in the regions of the Cygnus bubble and Geminga. The region of the Cygnus bubble is masked within about 6, while the region around Geminga had its contribution removed within an 8 radius, which is about 2.5 times larger than the masked regions of other sources listed in TeVCat. As shown in Figure 4, even with the removal of contributions from extended sources in a larger area, the Cygnus bubble and Geminga halo still provide dominant contributions. The detailed calculation process of the Cygnus bubble component in Figure 4 can be found in our previous work (Nie et al., 2024). Similar to the diffusion of electrons/positrons in the halo in this work, the Cygnus bubble is also considered a slow diffusion region, and its energy spectrum, morphological distribution, and diffuse gamma-ray structure in the region have been well reproduced.
@@ 3.2 Map of Geminga halo
In the vicinity of the Geminga pulsar, no radiation morphology was observed in the 1-8 GeV energy range, but there is extended radiation in the TeV range. To investigate the mechanism behind this phenomenon, we use the observational data of the multi-wavelength energy spectrum from the Geminga halo, as well as the model parameters derived from the constraints of the diffuse gamma-ray observations of the Galactic disk by LHAASO, to calculate the map signals of the Geminga halo and the background. As shown in Figure 5, in the 1-10 GeV energy range, the Geminga halo signal is relatively weak compared to the background intensity, Consequently, under background contamination, the extended emission morphology in the vicinity of the Geminga pulsar exhibits a low significance level. However, in the 1-100 TeV energy range, Geminga produces a signal much stronger than the background, and a significant radiation morphology remains even with the background contribution considered, consistent with the observations from the HAWC.
@@
Theoretical models suggest that lower-energy electrons/positrons produce the GeV radiation through inverse Compton scattering with soft background photons. However, once the electrons/positrons escape from the pulsar, they propagate through the slow diffusion region near it and diffuse into the more distant and homogeneous normal Galactic environment. In the region where ( ), the diffusion coefficient of the particles is at least two orders of magnitude smaller than in other regions due to the spatial-dependence of , meaning that the density of electrons/positrons that generate GeV radiation in this region is high. This situation appears inconsistent with the absence of significant extended emission detected within the region centered on Geminga (Xi et al., 2019). On the other hand, it has been suggested that the proper motion of Geminga becomes significant at GeV energies (Di Mauro et al., 2019). Therefore, the nondetection of few GeV emission can be explained by the scenario in which the region of interest ( ) considered is not large enough to detect the Geminga ICS halo due to the effect of Geminga’s proper motion.
we have incorporated and discussed the effects of the pulsar’s proper motion. The measured proper motion of the Geminga pulsar is in longitude and in latitude (Faherty et al., 2007). Assuming the current velocity is perpendicular to the line of sight, the pulsar’s spatial velocity vector is . We consider two scenarios: Firstly, the center of the Slow Diffusion Zone (SDZ) coincides with the Geminga pulsar, and the SDZ moves with the pulsar while its size scales proportionally with the square root of time. Secondly, the SDZ center remains at the pulsar’s birth location without co-motion. Normalization conditions are set as follows: For the first scenario, the SDZ size at the evolutionary endpoint (present time) is . For the second scenario, . The results are presented in Figure 6. It shows that for both scenarios (co-moving and static SDZ), pulsar proper motion significantly affects GeV-band radiation.
On the other hand, the diffuse gamma-ray “excess” has been observationally confirmed (Prodanović et al., 2007; Bartoli et al., 2015; Amenomori et al., 2021). We believe that signal leakage from sources like Geminga may provide a significant contribution to this phenomenon. Both scenarios demonstrate significant consistency with existing observational evidence: Fermi-LAT detected no significant 10–100 GeV emission within the region surrounding Geminga (Xi et al., 2019), while significant radiation is observed in the field at these energies (Di Mauro et al., 2019). We therefore calculate the contribution of Geminga halo signal leakage to the diffuse gamma-ray “excess” on the Galactic plane under this framework, as shown in the figure 7. As illustrated in the figure 7, multiplying Geminga’s leaked signal by a factor of 10 provides an excellent fit to the observed diffuse gamma-ray data. Here, we interpret this factor as indicative of multiple sources with similar leakage characteristics to Geminga distributed throughout this region.